OpenStudy (anonymous):

Please help.... Jennifer has 60 dollars more than Brian. If she were to give Brian 1/5 of her money, Brian would have 25% less than the amount that Jennifer would then have. How much money does Jennifer have?

OpenStudy (anonymous):

J=Jennifer amount of money B= Brian amount of money Jennifer has 60 dollars more than Brian: J=60+B "If she were to give Brian 1/5 of her money", then Brian has J/5+B, which is equal to "25% less than the amount that Jennifer would then have", or (4J/5)x75%. In mathematical terms, we can express the relationships with a linear system of two equations: J=60+B J/5+B=(4J/5)75% or J+5B=12J/4 or 4J+20B=12J. then B=2J/5 Therefore, J=60+2J/5 or J=100 B=40

OpenStudy (kobi):

100

OpenStudy (kobi):

hi